OpenStudy (anonymous):

Find the first four terms of the arithmetic sequence when a1 = 3 and d = 2. 2, 5, 8, 11 3, 5, 7, 9 1, 3, 5, 7 3, 6, 12, 24 PLEASE HEP :)

OpenStudy (anonymous):

Hello Kitty.. :)

OpenStudy (anonymous):

General Sequence of Arithmetic Progression is given by: \(a, a + d, a + 2d, a + 3d,.......................... a+ (n-1)d\)

OpenStudy (anonymous):

you are given \(a=3\) and \(d = 2\)... \(a\) is first term of the sequence and \(d\) is the common difference, it is called "Common", as difference between any two adjacent terms is equal.. :)

OpenStudy (anonymous):

First term \(= a = 3\) Second term \(=a +d = 3 + 2 = 5\) In other words,if you kee on adding \(2\) to the previous term, you will get next term.. :)

OpenStudy (anonymous):

*keep

OpenStudy (anonymous):

thank you so much :) @waterineyes